the sum of ages of two brothers is 24 in three years time, the elder brother will be twice as old as younger. how old are they ?
Answer by stanbon(75887)

the sum of ages of two brothers is 24.
a + b = 24
-----------------------------
in three years time, the elder brother will be twice as old as younger.
(a+3) = 2(b+3)
-----------------------------
how old are they ?
EQUATIONS Rearranged
a + b = 24
a -2b = 3
------------
Subtract the 2nd from the 1st to get:
3b = 21
b = 7 yrs. old (Younger brother's age)
------
Since a + b = 24 , a = 24-7 = 17 yrs. old (older brother's age)
=============================================
Cheers,
Stan H.